# Continual Learning for Multiple Modalities

### v1

@@ 3.3 Knowledge Aggregation
which guides the contribution of each prompt to the aggregated prompt ^$\bm{\hat{}}{P}^{t}_{m}$. Note that the text prompts do not require aggregation across modalities because text data is input solely into the language encoder. By aligning the modality and text embeddings extracted with the prompts, ^$\bm{\hat{}}{P}^{t}_{m}$ and ¯$\bm{\bar{}}{Q}^{t}_{m}$, we can avoid interference between modalities. To summarize, we update the set of parameters $\{P^{t}_{m},Q^{t}_{m},\alpha^{t}_{m}\}$ by minimizing the aforementioned loss functions.
@@
Re-alignment modality projection head. Learning new knowledge without accessing data from old tasks can cause the embeddings from the modality encoder to be biased toward the text embeddings of the latest task. To address this issue, we introduce an additional phase to re-align the projection head $f_{\bm{\bar{}}{\alpha}^{t}_{m}}(\cdot)$ so that old and new modality embeddings are distinguished. Specifically, for each class of modality $m$, we collect features mimicking old features, which are sampled using $V(\cdot;\bm{\hat{}}{P}^{i}_{m})$, where $i<t$. The aggregated projection head, ¯$\bm{\bar{}}{\alpha}^{t}_{m}$, is further fine-tuned to align the modality embedding derived from the sampled features with the text embedding by minimizing $\mathcal{L}_{m}$.
4 Experiments
4.1 Setup
Datasets. We evaluated the proposed method for continual learning of five modalities: image, video, audio, depth, and text. The datasets used for the modalities are ImageNet-100 [5], UCF-101 [36], ESC-50 [25], and SUN-RGBD [35], each accompanied by corresponding text data. ImageNet-100 contains 100 classes selected from the original ImageNet dataset [5], widely used in continual learning for image classification [7, 47]. UCF-101 contains 101 action categories for video action recognition [36]. ESC-50 is a fine-grained collection of environmental audio recordings of 50 classes for sound classification. For the depth modality, we performed on the SUN-RGBD dataset for scene classification. Each dataset was divided into five non-overlapping subsets of classes. We consider each subset as a task. Please refer to the supplementary material for preprocessing each modality dataset and the class orders.
Scenarios. We compared the proposed method with other methods with three sequences of tasks: (i) A random modality is introduced at each time step. (ii) A modality is trained across multiple time steps before shifting to the next modality. (iii) All modalities are available at every time step. The first two sequences comprise 20 time steps (corresponding to five subsets of four modalities), with each task exclusively learning a single modality. In contrast, the third sequence consists of five time steps, during which all modalities are learned simultaneously for every step. For each sequence, we evaluated the methods with and without the provision of modality identity during the evaluation phase, which we refer to as the modality-specific and modality-agnostic scenarios, respectively. We report average incremental accuracy (AIA) [30, 7], final average accuracy (FAA) [7, 39], and average forgetting (F) [3, 33].
Implementation details. For a fair comparison with other continual learning methods, we adopted ViT-B-16 CLIP [28] as the backbone model for the compared methods. We used the vision and text encoders of CLIP for the non-text and text modalities, respectively, in accordance with established practices [9, 53]. We used a fully connected layer for $g_{\beta^{t}}(\cdot)$. Additionally, due to the high dimension of the projection layer $\alpha^{t}_{m}$, we employed a low-rank decomposition technique [12] to split it into two smaller matrices with a rank of 20. The proposed method was trained using the Adam optimizer [15] with $\beta_{1}$ and $\beta_{2}$ of 0.9 and 0.999. We report the average results from three independent runs for all experiments using NVIDIA RTX A5000 GPUs. For additional details on the implementation, including the preprocessing strategies, please refer to the supplementary material.
Baselines. We evaluated the proposed method, COMM, with existing continual learning methods, including sequential fine-tuning (FT), EWC [16], LwF [17], WISE-FT [45], L2P [44], S-liPrompts [41], AttriCLIP [40], and PGP [27]. For the conventional methods, EWC, LwF, and WISE-FT, we directly updated the parameters of modality and text encoders. The prompt-based methods, L2P [^3] Since L2P is a ViT-based method [6], we incorporated the prompts into the modality encoder in the proposed framework for comparison., AttriCLIP, and PGP, used modality-specific and modality-shared prompt pools for the modality-specific and modality-agnostic class-incremental learning scenarios.
Table 1: Results for the first sequence on the modality-specific (top) and modality-agnostic (bottom) class-incremental learning scenarios. The AIA, FAA, and F metrics are measured after training all tasks. Overall denotes the average AIA and FAA across all modalities.
@@ 4.1 Setup
EWC [16] 47.16 23.42 24.98 48.59 20.76 29.37 26.38 13.91 13.20 26.68 5.50 20.50 37.20 15.90 22.51
LwF [17] 53.13 24.78 29.84 48.59 23.26 26.74 30.64 11.43 20.34 31.89 9.75 23.53 41.06 17.31 25.11
WISE-FT [45] 46.73 21.96 26.08 46.99 21.27 27.14 26.93 5.27 22.93 22.69 8.25 15.34 35.83 14.18 22.87
L2P [44] 76.96 68.66 8.74 78.49 70.71 8.21 45.06 24.75 24.66 44.28 17.00 28.98 61.20 45.28 17.65
S-liPrompts [41] 75.61 65.25 10.91 82.85 69.04 14.57 47.98 26.08 30.07 42.68 22.25 21.70 62.28 45.66 19.31
AttriCLIP [40] 77.39 65.70 12.30 77.40 70.51 7.28 51.21 35.01 17.16 43.44 20.50 24.37 62.36 47.93 15.27
PGP [27] 78.88 68.58 10.84 83.14 71.46 12.33 48.67 25.58 24.45 47.69 26.25 22.78 64.60 47.98 17.60
@@
EWC [16] 38.23 5.60 34.34 27.46 3.63 25.15 24.04 13.91 10.73 22.77 5.50 18.34 28.13 7.15 22.14
LwF [17] 48.74 21.14 29.05 19.82 1.17 22.43 23.01 6.74 17.22 19.01 9.00 10.70 27.65 9.51 19.85
WISE-FT [45] 38.43 10.34 29.57 24.89 0.36 25.88 25.94 5.27 21.67 22.46 6.25 17.22 27.93 5.55 23.59
L2P [44] 79.85 69.84 10.53 71.25 65.14 6.44 39.92 26.72 13.96 33.08 21.50 12.31 56.03 45.80 19.31
S-liPrompts [41] 77.36 66.72 11.20 79.52 73.18 6.69 40.94 30.02 11.56 33.63 23.50 10.79 57.87 48.36 10.06
AttriCLIP [40] 73.93 63.90 10.56 73.28 68.70 4.83 39.52 28.57 11.60 34.99 22.50 13.27 55.43 45.91 10.06
PGP [27] 81.29 70.92 10.85 77.55 67.18 10.94 45.61 26.25 20.49 46.15 21.50 26.19 62.37 41.58 17.12
@@
4.2 Main Results
Sequence 1. We evaluated our method and the compared methods on the first sequence for multimodal continual learning. Table 1 summarizes the results of the modality-specific (top) and modality-agnostic (bottom) class-incremental learning scenarios. Overall, we observe that the results of the modality-agnostic scenario show lower performance compared to the modality-specific scenario due to the challenge in identifying the modality. The regularization methods, EWC, LwF, and WISE-FT, perform poorly in both scenarios compared to the prompt-based continual learning methods, L2P, S-liPrompts, AttriCLIP, PGP, and COMM. This indicates that directly updating the network parameters in the regularization methods to accommodate knowledge from multiple modalities is prone to forgetting previously learned tasks due to contaminated mixed knowledge. Even though the existing prompt-based methods outperform other continual learning approaches, they do not perform well for the depth and audio modalities. Notably, these methods exhibit lower AIA in the modality-agnostic scenario due to the challenge of selecting prompts from the shared pool, which hinders learning new tasks. The proposed method, aiming to alleviate sensitivity to prompt selection and biased projections, outperforms the best competitor, S-liPrompts, with an overall FAA gap of 16.85% and 13.56% in the respective scenarios. In terms of the performance gap between the scenarios, the proposed method shows only a marginal performance gap of 0.63% in overall AIA, in contrast to the larger gaps of 4.41% and 6.93% observed with S-liPrompts and AttriCLIP, respectively. Additional results for different class orders can be seen in the supplementary material.
@@ 4.2 Main Results
EWC [16] 27.21 28.45 18.19 30.32 26.04 24.86 26.10 24.37 16.98 29.88 24.33 26.59
LwF [17] 26.92 28.52 18.71 27.76 25.48 24.83 20.48 23.12 18.23 24.45 21.57 21.49
WISE-FT [45] 27.08 27.85 16.53 29.17 25.16 24.67 23.43 24.00 16.01 28.02 22.87 24.63
L2P [44] 70.73 72.96 32.25 39.54 53.87 11.25 72.19 62.37 37.03 43.38 53.74 11.95
S-liPrompts [41] 67.93 75.45 40.88 43.25 56.88 12.02 67.81 73.42 32.17 38.62 53.01 11.29
AttriCLIP [40] 68.62 72.69 42.30 42.24 56.46 10.43 67.89 69.91 38.26 44.05 55.03 12.62
PGP [27] 72.51 76.51 37.42 43.87 57.58 10.11 72.15 66.24 37.22 42.72 54.58 11.31
@@
AGG SR PHR Image Video Depth Audio Overall Image Video Depth Audio Overall
@@
Table 4: Results of the ablation study on prompting designs.
Method Image Video Depth Audio Overall
COMM-L 85.30 83.89 61.56 60.43 72.80
COMM-V 85.76 90.50 59.22 61.35 74.21
COMM-VL 86.07 91.03 62.90 61.86 75.47
@@
4.3 Ablation Study and Analysis
Ablation study on the components. We conducted an ablation study of the proposed method, including its components: prompt aggregation (AGG), self-regularization (SR), and projection head re-alignment (PHR). Note that the method without AGG or SR requires identifying the modality identity or task to select prompts, respectively. Therefore, we employed a prompt selection mechanism in [41]. We compared these methods to the first two sequences in the modality-agnostic scenario. Table 3 reports the results of the ablation study with AIA measured after training all tasks. Overall, the proposed method using an individual component results in unsatisfactory performance on both task sequences compared to employing them together. The method using AGG identifies relevant modalities but struggles to select accurate task-specific prompts. When comparing the method utilizing all components against that excluding AGG, it becomes evident that selecting suitable modality prompts significantly enhances performance, with gaps of 15.23% and 20.38% in the depth modality for the first and second sequences, respectively. The method that employs both SR and PHR exhibits a strong synergistic effect because SR preserves previous features by regularizing prompts, and these preserved features closely resemble those used by PHR. We provide visualizations demonstrating the maintainability of SR in the embedding space in the supplementary material.
@@ 4.3 Ablation Study and Analysis
Ablation study on the prompting designs. We conducted an additional study to investigate the impact of prompting design strategies in continual learning. In this study, we compared the proposed method, which uses prompts for both modality and language encoders (which we call COMM-VL), with the method that uses prompts for the modality (COMM-V) or the language encoder (COMM-L). We applied these methods to the first sequence in the modality-specific scenario. In Table 4, we observe that COMM-VL yields higher overall AIA performance than COMM-V and COMM-L. COMM-L gives the lowest performance among the compared approaches due to its insufficiency in learning diverse modalities, highlighting the requirements for prompts in the modality encoder.
Study of scalability. In this study, we evaluated the scalability of the proposed method by examining the impact of varying numbers of learnable parameters. To adjust the number of learnable parameters, we increased the dimension of prompts and the layers of the pre-trained models for inserting prompts, following the methods in the previous studies [14, 51]. Figure 5 indicates that COMM maintains consistent overall AIA performance across different prompt dimensions and numbers of layers for prompt insertion. Furthermore, we also compared the parameter growth of COMM with other prompt-based methods in Figure 6. Overall, prompt-based methods require fewer parameters than updating the entire model [16, 17], which involves approximately 124 million parameters. Specifically, L2P and AttriCLIP leverage a prompt pool for each modality, resulting in a step-wise increase in parameters as new modalities are introduced. S-liPrompts uses task-specific prompts, leading to a monotonic parameter increase. The proposed method, COMM, consolidates prompts inter- and intra-modalities, enhancing scalability when integrating additional modalities.

### v2

@@ 3.3 Continual Learning for Multiple Modalities
This composed prompt $\bm{\hat{}}{P}^{t}_{m}$ selectively emphasizes prompts associated with modality contexts relevant to the current input, while attenuating the influence of unrelated ones. Note that the text prompts do not require such aggregation because text data is input solely into the language encoder. By aligning the modality and text embeddings extracted using the prompts, $\bm{\hat{}}{P}^{t}_{m}$ and $\bm{\bar{}}{Q}^{t}_{m}$, we can avoid interference between modalities. To summarize, we update the set of parameters $\{P^{t}_{m},Q^{t}_{m},\alpha^{t}_{m}\}$ by minimizing the aforementioned loss functions.
@@
Modality-text re-alignment. Learning new knowledge without accessing data from old tasks can cause the embeddings from the modality encoder to be biased toward the text embeddings of the latest task (Yu et al. 2020), which distorts the previously modality-text alignments. This degrades the semantic grounding of accumulated prompts, as they are composed based on earlier alignments, making the modality-text inconsistency more severe. To address this issue, we introduce an additional phase to re-align the projection head, $f_{\bm{\bar{}}{\alpha}^{t}_{m}}(\cdot)$, so that old and new modality embeddings are distinguished. Specifically, for each class of modality $m$, we sample features that mimic old ones using the class-wise mean and covariance of previously stored embeddings, which are extracted via the modality encoder with the prompt, $\bm{\hat{}}{P}^{i}_{m}$, where $i<t$. The aggregated projection head, $\bm{\bar{}}{\alpha}^{t}_{m}$, is further fine-tuned to align the modality embedding derived from the sampled features with the text embedding by minimizing $\mathcal{L}_{\text{task}}$.
@@
EWC (Kirkpatrick et al. 2017) 47.16 23.42 24.98 48.59 20.76 29.37 26.38 13.91 13.20 26.68 5.50 20.50 37.20 15.90 22.51
LwF (Li and Hoiem 2016) 53.13 24.78 29.84 48.59 23.26 26.74 30.64 11.43 20.34 31.89 9.75 23.53 41.06 17.31 25.11
WISE-FT (Wortsman et al. 2022) 46.73 21.96 26.08 46.99 21.27 27.14 26.93 5.27 22.93 22.69 8.25 15.34 35.83 14.18 22.87
L2P (Wang et al. 2022) 76.96 68.66 8.74 78.49 70.71 8.21 45.06 24.75 24.66 44.28 17.00 28.98 61.20 45.28 17.65
S-liPrompts (Wang, Huang, and Hong 2022) 75.61 65.25 10.91 82.85 69.04 14.57 47.98 26.08 30.07 42.68 22.25 21.70 62.28 45.66 19.31
AttriCLIP (Wang et al. 2023) 77.39 65.70 12.30 77.40 70.51 7.28 51.21 35.01 17.16 43.44 20.50 24.37 62.36 47.93 15.27
PGP (Qiao et al. 2024) 78.88 68.58 10.84 83.14 71.46 12.33 48.67 25.58 24.45 47.69 26.25 22.78 64.60 47.98 17.60
@@
EWC (Kirkpatrick et al. 2017) 38.23 5.60 34.34 27.46 3.63 25.15 24.04 13.91 10.73 22.77 5.50 18.34 28.13 7.15 22.14
LwF (Li and Hoiem 2016) 48.74 21.14 29.05 19.82 1.17 22.43 23.01 6.74 17.22 19.01 9.00 10.70 27.65 9.51 19.85
WISE-FT (Wortsman et al. 2022) 38.43 10.34 29.57 24.89 0.36 25.88 25.94 5.27 21.67 22.46 6.25 17.22 27.93 5.55 23.59
L2P (Wang et al. 2022) 79.85 69.84 10.53 71.25 65.14 6.44 39.92 26.72 13.96 33.08 21.50 12.31 56.03 45.80 19.31
S-liPrompts (Wang, Huang, and Hong 2022) 77.36 66.72 11.20 79.52 73.18 6.69 40.94 30.02 11.56 33.63 23.50 10.79 57.87 48.36 10.06
AttriCLIP (Wang et al. 2023) 73.93 63.90 10.56 73.28 68.70 4.83 39.52 28.57 11.60 34.99 22.50 13.27 55.43 45.91 10.06
PGP (Qiao et al. 2024) 81.29 70.92 10.85 77.55 67.18 10.94 45.61 26.25 20.49 46.15 21.50 26.19 62.37 41.58 17.12
@@
Table 1: Results for the first sequence on the modality-specific (top) and modality-agnostic (bottom) scenarios. The AIA, FAA, and F metrics are measured after training all tasks. Overall denotes the average AIA and FAA across all modalities.
@@
4 Experiments
4.1 Experimental Setup
Scenarios. We evaluated the proposed method for continual learning of five modalities: image, video, audio, depth, and text, using four datasets: ImageNet-100 (Deng et al. 2009), UCF-101 (Soomro, Zamir, and Shah 2012), ESC-50 (Piczak 2015), and SUN-RGBD (Song, Lichtenberg, and Xiao 2015). Each dataset was divided into five non-overlapping subsets of classes. We consider each subset as a task. For a comprehensive evaluation, we designed a benchmark based on three sequences of tasks: (i) A random modality is introduced at each time step. (ii) A modality is trained across multiple time steps before shifting to the next modality. (iii) All modalities are available at every time step. The first two sequences comprise 20 time steps (corresponding to five subsets of four modalities), with each task exclusively learning a single modality. In contrast, the third sequence consists of five time steps, during which all modalities are learned simultaneously for every step. For each sequence, we evaluated the methods with and without the provision of modality identity during the evaluation phase, which we refer to as the modality-specific and modality-agnostic scenarios, respectively. We report average incremental accuracy (AIA) (Rebuffi et al. 2017), final average accuracy (FAA) (Douillard et al. 2022), and average forgetting (F) (Chaudhry et al. 2019).
Baselines. We evaluated the proposed method, COMM, with existing continual learning methods: conventional approaches, including sequential fine-tuning (FT), EWC (Kirkpatrick et al. 2017), LwF (Li and Hoiem 2016), and WISE-FT (Wortsman et al. 2022); and prompt-based methods, L2P (Wang et al. 2022), S-liPrompts (Wang, Huang, and Hong 2022), AttriCLIP (Wang et al. 2023), and PGP (Qiao et al. 2024).
Implementation details. For a fair comparison with other continual learning methods, we adopted ViT-B-16 CLIP (Radford et al. 2021) as the backbone model for the compared methods. We used the vision and text encoders of CLIP for the non-text and text modalities, respectively, in accordance with established practices (Girdhar et al. 2023; Zhu et al. 2023). We used a fully connected layer for $g_{\beta^{t}}(\cdot)$. The proposed method was trained using the Adam optimizer (Kingma and Ba 2015) with $\beta_{1}$ and $\beta_{2}$ of 0.9 and 0.999. Additional information on datasets, preprocessing per modality, baseline setups, and hyperparameters are provided in the supplementary material.
@@ 4.1 Experimental Setup
EWC (Kirkpatrick et al. 2017) 27.21 28.45 18.19 30.32 26.04 24.86 26.10 24.37 16.98 29.88 24.33 26.59
LwF (Li and Hoiem 2016) 26.92 28.52 18.71 27.76 25.48 24.83 20.48 23.12 18.23 24.45 21.57 21.49
WISE-FT (Wortsman et al. 2022) 27.08 27.85 16.53 29.17 25.16 24.67 23.43 24.00 16.01 28.02 22.87 24.63
L2P (Wang et al. 2022) 70.73 72.96 32.25 39.54 53.87 11.25 72.19 62.37 37.03 43.38 53.74 11.95
S-liPrompts (Wang, Huang, and Hong 2022) 67.93 75.45 40.88 43.25 56.88 12.02 67.81 73.42 32.17 38.62 53.01 11.29
AttriCLIP (Wang et al. 2023) 68.62 72.69 42.30 42.24 56.46 10.43 67.89 69.91 38.26 44.05 55.03 12.62
PGP (Qiao et al. 2024) 72.51 76.51 37.42 43.87 57.58 10.11 72.15 66.24 37.22 42.72 54.58 11.31
@@
4.2 Main Results
Sequence 1. We evaluated our method and the compared methods on the first sequence for multimodal continual learning. Table 1 summarizes the results of the modality-specific (top) and modality-agnostic (bottom) class-incremental learning scenarios. Overall, we observe that the results of the modality-agnostic scenario show lower performance compared to the modality-specific scenario due to the challenge in identifying the modality. The regularization methods, EWC, LwF, and WISE-FT, perform poorly in both scenarios compared to the prompt-based continual learning methods, L2P, S-liPrompts, AttriCLIP, PGP, and COMM. This indicates that directly updating the network parameters in the regularization methods to accommodate knowledge from multiple modalities is prone to forgetting previously learned tasks due to contaminated mixed knowledge. Even though the existing prompt-based methods outperform other continual learning approaches, they do not perform well for the depth and audio modalities. This is mainly due to the absence of pre-training on non-visual data, which leads to greater forgetting; the proposed method alleviates this by effectively weighting prompts and suppressing cross-modal interference. Notably, these methods exhibit lower AIA in the modality-agnostic scenario due to the challenge of selecting prompts from the shared pool, which hinders learning new tasks. The proposed method, aiming to alleviate sensitivity to prompt selection and biased projections, outperforms the best competitor, S-liPrompts, with an overall FAA gap of 16.85% and 13.56% in the respective scenarios. In terms of the performance gap between the scenarios, the proposed method shows only a marginal performance gap of 0.63% in overall AIA, in contrast to the larger gaps of 4.41% and 6.93% observed with S-liPrompts and AttriCLIP, respectively. Additional results for different class orders can be seen in the supplementary material.
@@ 4.2 Main Results
Cross Self Re-align Image Video Depth Audio Overall Image Video Depth Audio Overall
@@
4.3 More Results
Ablation study on the components. We conducted an ablation study of the proposed method, including its components: cross-modality integration (Cross), self-regularization (Self), and projection head re-alignment (Re-align). Note that the method without Cross or Self requires identifying the modality identity or task to select prompts, respectively. Therefore, we employed a prompt selection mechanism in (Wang, Huang, and Hong 2022). We compared these methods to the first two sequences in the modality-agnostic scenario. Table 3 reports the results of the ablation study with AIA measured after training all tasks. Overall, the proposed method using an individual component results in unsatisfactory performance on both task sequences compared to employing them together. The method using Cross identifies relevant modalities but struggles to select accurate task-specific prompts. When comparing the method utilizing all components against that excluding Cross, it becomes evident that selecting suitable modality prompts significantly enhances performance, with gaps of 15.23% and 20.38% in the depth modality for the first and second sequences, respectively. The method that employs both Self and Re-align exhibits a strong synergistic effect because Self preserves previous features by regularizing prompts, and these preserved features closely resemble those used by Re-align. We provide visualizations demonstrating the maintainability of Self in the embedding space in the supplementary material.
@@ 4.3 More Results
Study of scalability. In this study, we evaluated the scalability of the proposed method by examining the impact of varying numbers of learnable parameters. To adjust the number of learnable parameters, we increased the dimension of prompts and the layers of the pre-trained models for inserting prompts, following the methods in the previous studies (Khattak et al. 2023; Zhou et al. 2022a). Figure 5 indicates that COMM maintains consistent overall AIA performance across different prompt dimensions and numbers of layers for prompt insertion. Furthermore, we also compared the parameter growth of COMM with other prompt-based methods in Figure 6. Overall, prompt-based methods require fewer parameters than updating the entire model (Kirkpatrick et al. 2017; Li and Hoiem 2016), which involves approximately 124 million parameters. Specifically, L2P and AttriCLIP leverage a prompt pool for each modality, resulting in a step-wise increase in parameters as new modalities are introduced. S-liPrompts uses task-specific prompts, leading to a monotonic parameter increase. The proposed method, COMM, consolidates prompts inter- and intra-modalities, enhancing scalability when integrating additional modalities. Additional results on prompting designs, prompt accumulation strategies, and unimodal continual learning are included in the supplementary material.